Calculate the total rates of return (holding period returns) for the following stocks:

Stock Beginning price: P0 Ending price: P1 Dividend payment: D1
A 28.38 23.06 2.6
B 110.74 120.92 5.25
C 4.25 5.58 0.24
What was the rate of return for stock A?

What was the rate of return for stock B?

What was the rate of return for stock C?

2. Calculate the total rates of return (holding period returns) for the following stocks: Stock Beginning price: Po Ending price: P, Dividend payment D 28.38 110.74 4.25 23.06 120.92 5.58 2.6 5.25 0.24 a. What was the rate of return for stock A? b. What was the rate of return for stock B? c. What was the rate of return for stock C?
Answer #1

Solution-

Holding period return = [( Ending price{P1} - Beginning price {P0})+Dividend {D1}]/Beginning Price {P0}

Stock A:-

Beginning Price {P0}= $28.38

Ending Price {P1}= $23.06

Dividend {D1)= $2.6

Holding period return = (23.06-28.38)+2.6/28.38

Holding period return = -2.72/28.38= -0.0958 or -9.58%

Stock B:

Beginning price (P0) = $110.74

Ending price (P1)= $120.92

Dividend (D1)= $5.25

Holding period return = (120.92-110.74)+5.25/110.74

Holding period return= 15.43/110.74= 0.1393 or 13.93%

Stock C :

Beginning price (P0)= $4.25

Ending Price (P1)= $5.58

Dividend (D1) = $0.24

Holding period return = (5.58-4.25)+0.24/4.25

Holding period return= 1.57/4.25= 0.3694 or 36.94%